# On the spatial distribution of the Large-Scale structure: An Unsupervised search for Parity Violation

### v2

@@ Abstract
We use machine learning methods to search for parity violations in the Large-Scale Structure (LSS) of the Universe, motivated by recent claims of chirality detection using the 4-Point Correlation Function (4PCF), which would suggest new physics during the epoch of inflation. This work seeks to reproduce these claims using methods originating from high energy collider analyses. Our machine learning methods optimise some underlying parity odd function of the data, and use it to evaluate the parity odd fraction. We demonstrate the effectiveness and suitability of these methods and then apply them to the Baryon Oscillation Spectroscopic Survey (BOSS) catalogue. No parity violation is detected at any significance.
@@ IV Validating the Methods
To begin, in Section IV.1 we use the CNN architecture of Lester and Tombs [37] to examine a two-dimensional projection of the data. This is followed Section IV.2 by the addition of colour channels to the images, with the aim of representing three-dimensional information via colour. Finally, we investigate more geometric approaches in Section IV.3, which have the drawback of being constrained to some order of polyhedron, much like $N$-point correlation functions, but consequently allow direct comparison with the 4PCF.
@@ V.4 Results
The average level of detection from bootstrapping over the testing sets and their controls for each method, as well as the random toy tests for comparison are given in Table 7. For every method, no significant parity violation was detected. This is confirmed in the $\sigma$ levels displayed in Table 8. These show no significant deviation from the control groups or from random generated data.
@@
2D Images $50.4\pm 0.3\%$ $50.1\pm 0.3\%$ $49.9\pm 0.3\%$
Colour Images $49.7\pm 0.3\%$ $49.4\pm 0.3$ $50.2\pm 0.3\%$
2D Angles, n=3 $50.15\pm 0.09\%$ $50.02\pm 0.09\%$ $50.41\pm 0.10\%$
2D Angles, n=4 $50.05\pm 0.06\%$ $49.85\pm 0.06\%$ $50.16\pm 0.05\%$
2D Angles, n=5 $50.30\pm 0.06\%$ $50.27\pm 0.06\%$ $49.60\pm 0.05\%$
3D Angles $49.91\pm 0.06\%$ $50.09\pm 0.05\%$ $49.92\pm 0.05\%$
3D Vectors $50.08\pm 0.03\%$ $50.09\pm 0.03\%$ $50.03\pm 0.03\%$
Table 7: The average deviation from parity even of the CMASS NGC catalogue for the different methods.
@@
2D Images 0.662 0.662
Colour Images 1.00 0.333
2D Angles, n=3 1.02 1.93
2D Angles, n=4 1.17 1.41
2D Angles, n=5 0.353 1.28
3D Angles 0.00 0.128
3D Vectors 0.235 1.18
Table 8: The $\sigma$-level deviation, assuming Gaussian error, calculated from differences from 50%. This shows no significant parity violations.
@@ VI Discussion
The final important consideration is the ‘artifacts’ of the dataset. The 2d projection in Fig. 2 displays notable sample artifacts, which could interfere with the analyses. In all methods, the weights applied Section III act to account for sampling issues, reducing the impact. In the case of the CNN, images were constructed to further offset the impact: sampling a large enough set of small enough patches minimised overlap with artifact regions, displayed in Fig. 13. That the overall impact is negligible is perhaps best displayed by the null results in all cases. Given that no parity violation is detected, if the artifacts were interfering, it would have to be in a way that perfectly cancels with some underlying parity violation of the dataset. Since no detection is made in either 2d or 3d, the probability of this being the case is negligible. To make one final confirmation of this belief, the sky was split into 8 bins by position, displayed in Fig. 14. Treating each of these regions separately, no significant difference was found between regions (Appendix D). Since every region is differently affected by artifacts, this suggests they have no noticeable impact on results.
@@
Figure 13: The set of images generated from points near the large artifact between (210, 35) and (240, 50). Very few images are affected by the presence of the artifact, and those which are affected have small overlaps into the affected region.
@@ Acknowledgements.
SH thanks the Cavendish Laboratory for the Part III Project opportunity. SH also thanks St John’s College and the University of Cambridge for funding throughout the year. WH was supported by a Royal Society University Research Fellowship.

### v3

@@ Abstract
We use machine learning methods to search for parity violations in the Large-Scale Structure (LSS) of the Universe, motivated by recent claims of chirality detection using the 4-Point Correlation Function (4PCF), which would suggest new physics during the epoch of inflation. This work seeks to reproduce these claims using methods originating from high energy collider analyses. Our machine learning methods optimise some underlying parity odd function of the data, and use it to evaluate the parity odd fraction. We demonstrate the effectiveness and suitability of these methods and then apply them to the Baryon Oscillation Spectroscopic Survey (BOSS) catalogue. No strong evidence for parity violation is detected.
@@ IV Validating the Methods
To begin, in Section IV.1 we use the CNN architecture of Lester and Tombs [37] to examine a two-dimensional projection of the data. This is followed in Section IV.2 by the addition of colour channels to the images, with the aim of representing three-dimensional information via colour. Finally, we investigate more geometric approaches in Section IV.3, which have the drawback of being constrained to some order of polyhedron, much like $N$-point correlation functions, but consequently allow direct comparison with the 4PCF.
@@ V.4 Results
The average level of detection from bootstrapping over the testing sets and their controls for each method are given in Table 7. We also include results from running over the random catalogue provided by BOSS. For every method, no significant deviation from the control or random data is detected. This is confirmed in the $\sigma$ levels displayed in Table 8. We do not find evidence of parity violation.
@@
2D Images $50.4\pm 0.3\%$ $50.1\pm 0.3\%$ $50.8\pm 0.3\%$
Colour Images $49.7\pm 0.3\%$ $49.4\pm 0.3\%$ $49.3\pm 0.3\%$
2D Angles, n=3 $50.15\pm 0.09\%$ $50.02\pm 0.09\%$ $50.35\pm 0.10\%$
2D Angles, n=4 $50.05\pm 0.06\%$ $49.85\pm 0.06\%$ $50.20\pm 0.06\%$
2D Angles, n=5 $50.30\pm 0.06\%$ $50.27\pm 0.06\%$ $49.65\pm 0.06\%$
3D Angles $49.91\pm 0.06\%$ $50.09\pm 0.05\%$ $49.84\pm 0.05\%$
3D Vectors $50.08\pm 0.03\%$ $50.09\pm 0.03\%$ $50.13\pm 0.03\%$
Table 7: The average deviation from parity even of the CMASS NGC catalogue for the different methods. Results for the control groups and random catalogue are also shown.
@@
2D Images 1.00 1.33
Colour Images 1.00 1.33
2D Angles, n=3 1.44 2.00
2D Angles, n=4 1.67 2.50
2D Angles, n=5 0.500 0.833
3D Angles 0.00 1.00
3D Vectors 0.333 1.67
Table 8: $\sigma$-level significances calculated for both (i) the difference between the CMASS data and the CMASS controls, and (ii) the difference between the CMASS data and the BOSS random catalogue data. No significant parity violation is found.
@@ VI Discussion
An important consideration is the ‘artifacts’ of the dataset. The 2d projection in Fig. 2 displays notable sample artifacts, which could interfere with the analyses. In all methods, the weights applied Section III act to account for sampling issues, reducing the impact. For the CNN, images were constructed to further offset the impact: sampling a large enough set of small enough patches minimised overlap with artifact regions, displayed in Fig. 13. That the overall impact is negligible is perhaps best displayed by the null results in all cases. Given that no parity violation is detected, if the artifacts were interfering, it would have to be in a way that perfectly cancels with some underlying parity violation of the dataset. Since no detection is made in either 2d or 3d, the probability of this being the case is negligible. To further the rigorousity of these claims, the analyses were performed over the random catalogue of the BOSS dataset. This catalogue is constructed to match the sampling geometry of the BOSS data, but is $\sim$50 times the size. To resolve this size mismatch, the random data was split into 50 chunks, 10 of which were used for the analysis. The results given in Table 8 show that sampling geometry indeed introduces no parity violation. To make a final confirmation, the sky was split into 8 bins by position, displayed in Fig. 14, and inspired by Hou et al. [17]. Treating each of these separately, no significant difference was found between regions (Appendix D). Since every region is differently affected by artifacts, this suggests they have no noticeable impact on results.
A final important consideration is to note that the errors quoted for each dataset represent the model variance arising from our network and the resampling process. Cosmic variance from the single realisation of the Universe will contribute an extra source of uncertainty to the BOSS catalogue, such that even in the limit of infinite samples of the data, some underlying non-zero baseline variance would remain.
@@
Figure 13: The set of images generated from points near the large artifact between (210, 35) and (240, 50). Few images are affected by the presence of the artifact, and those which are affected have small overlaps into the affected region.
@@ Acknowledgements.
SH thanks the Cavendish Laboratory for the Project opportunity. SH also thanks St John’s College and the University of Cambridge for funding. WH was supported by a Royal Society University Research Fellowship.
VIII Data Availability
The data that support the findings of this article are openly available [52, 53].